OpenStudy (muddahamburglar):

f(x)= 3tan(2x) i think the answer is pi/2 but I'm not sure.

OpenStudy (priyar):

what is the q?

OpenStudy (salazarblack):

what do you need to do with this?

OpenStudy (muddahamburglar):

oh the period of the function

OpenStudy (salazarblack):

it is pi/2 * n

OpenStudy (muddahamburglar):

or would it be 2pi bc i need to multiply pi/2 by the coefficient of x?

OpenStudy (salazarblack):

As a general formula, "a" shows the amplitude of the function. b can be used to calculate the period of the function, where period = π/b (for a tan function) Thus amplitude = a = 3 period = π/b = π/2

OpenStudy (muddahamburglar):

so it is pi/2 ?

OpenStudy (salazarblack):

yup